today's access

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Wed Aug 29 2001 - 21:42:39 EDT

  • Next message: Apache: "Shift report 20010830 16:00-24:00"

    Yesterday I negotitiated an access in order to a)fix the HV problem and b) t T3 out of IR in order for Zbigniew to
    perform necessary repairs. This morning a second wire had broken in the 2. module so this repair is timely.
    The activities were as I had reported to the scheduling physicist.
    
    a) The T3 DC was removed and the second module (of 3) which had fail moved
    out of the IR for repair. We should definitely be ready
    by next Thursday for re-installation, but it is quite likely we will be
    ready by the weekend, and would be able to install should there be an
    unscheduled machine downtime. (and riggers available)
    
    b) We do not know with 100% certainty that the HV problem was fixed, but 2
    of the 6 modules in the chain showed some errors
    during the testing in the afternoon opening, and was replaced. At the end of
    the day communication was re-established to all modules (main frames) and
    hopefully this is now ok.
    
    I should point out that RHIC is planning to run over labour day weekend (wed too is a lab holiday) in production mode, thus making it unlikely we can re-install T3 until later next week.
    
    I will like to thank Bob, Erik, Alexandru, Hiro and Zbig for most valuable effort to these repairs.
    
    Dana, Andrev and myself has worked on the ZDC as announched yesterday, and they should be close to a good
    operation condition. We still have to do final checks and adjustments.
    The beam contiinues to have increased intensity, and fairly shortly it will become necessary to implement higfher order triggers
    both the centrality (in place) and a vertex cut (still to be done).
    
    As a last point the upcoming schedule is based on one-person shifts. It thus becomes imperative that shifters are well aware what to do 
    i.e. checking scalers, spectra, checking RHIC conditions. I do honestly not expect that when beam is available that shifters will have much time
    for doing other actrivities on shift. Please help to make the run a succesfull one, report problems and progress in the shift logs.
    
    Flemming
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Wed Aug 29 2001 - 21:34:07 EDT